OpenStudy (irish_waffl3):

Can some one help me factor 3x^2+5x+1

OpenStudy (irish_waffl3):

A. (3x+1)(x+1) B. (3x+5)(x+1) C. (3x-5)(x+1) D. Prime

OpenStudy (anonymous):

this cant be factored

OpenStudy (anonymous):

in other words D

OpenStudy (irish_waffl3):

Thank you!

OpenStudy (irish_waffl3):

Can i ask another one?

OpenStudy (anonymous):

Trinomial can not be factored

OpenStudy (trojanpoem):

maybe try. let it = 0 just for ease ( will omit that later) 3x^2+5x+1 = 0 3x^2 + 5x = -1 x^2 + 5/3 x = - 1/3 (x+5/6)^2 - 25/36 = -1/3 (x+5/6)^2 = 13/36 x + 5/6 = + - sqrt(13)/6 x = [-5 +- sqrt(13) ] /6 so ( x - ([-5 + sqrt(13)]/6))( x + ([-5+sqrt(13]/6))